A Young’s double-slit apparatus is set up so that a screen is positioned 2.1 m from the double slits, and the spacing between the two slits is 0.033 mm. What is the distance between alternating bright fringes on the screen if the light source has a wavelength of 678 nm? (1 nm = 10 −9 m)
Answer #1

D=201m d = 0.033 mm a = 678 nm xn= DAD xn+1 = (n+1) 2D d xnti - x = (n+1)D -nad = 678*10 x 2.1 0.033x 103 = 0.0431 m = 4. 31c

Please comment for any doubts.